the product of two consecutive positive odd integers is one less than three times their sum. find the integers.
Answer by mananth(16946)

let the number be x
the second number will be x+2
the product will be x(x+2)
three times the sum less 1 = 3(x+x+2)-1
x(x+2)=3(2x+2)-1
x^2+2x=6x+6-1
x^2+2x-6x=+6-1
x^2-4x=+5
x^2-4x-5=0
x^2-5x+x-5=0
x(x-5)+1(x-5)=0
(x+1)(x-5)=0
x=-1 or 5
So 5 & 7 are the numbers